Ophel has Php 53.00 in 5 peso and 1 peso coins pieces. If she has 25 coins all, how many has she of each kind?
Found 2 solutions by Cromlix, ikleyn:
Answer by Cromlix(4381)

Hi there
Number of coins = 25 coins
Type of coins = 53 Php
x + y = 25.......(1)
5x + y = 53......(2)
take (1) from (2)
4x = 28
x = 7
Substitute x = 7 in (1)
x + y = 25
7 + y = 25
y = 25 - 7
y = 18
7 x 5 peso coins and 18 x 1 peso coins
Hope this helps. :-)
Merry Christmas!

Answer by ikleyn(52788)

.

Worded by the same words and printed by the same symbols, it was solved a week or two weeks ago under this link

https://www.algebra.com/algebra/homework/coordinate/Linear-systems.faq.question.1150140.html

https://www.algebra.com/algebra/homework/coordinate/Linear-systems.faq.question.1150140.html

by exactly the same method with the same answer.